# LACOSTE: Exploiting stereo and temporal contexts for surgical instrument segmentation

### v1

@@ 1 Introduction
For SIS, the effectiveness of query-based segmentation (QBS) methods has been validated by previous works. We derive our method, which follows the QBS paradigm, grounded on three limitations of current methods for a surgical scenario. Firstly, surgical frames are recorded by stereo cameras and have inherent stereo attributes that are largely ignored by previous works. The complementary depth-aware information from stereo frame can enhance current view features, which help instrument localization and recognition for complex scenes. From this, we design a disparity-guided feature propagation (DFP) module, which integrates features from stereo views with an offline disparity estimation network explicitly. DFP can be inserted into QBS baseline without additional trainable parameters. To generalize well to a monocular setting, we also propose a pseudo stereo mechanism, which generates complementary right frames within multiple disparity ranges. Some examples are illstrated in Figure 3.
@@ 2 Related Work
$m$$\{m_{n}\}_{n=1}^{N}$ binary mask $p^{b}$$\{p^{b}_{n}\}_{n=1}^{N}$ class probability Final Output
$e$$\{e_{n}\}_{n=1}^{N}$ object query $m^{b}$$\{m^{b}_{n}\}_{n=1}^{N}$ binary mask $p^{f}$$\{p^{f}_{n}\}_{n=1}^{N}$ class probability
$\mathcal{Q}$ $\mathcal{Q}=\{q_{n}\}_{n=1}^{N}$ initial query $e^{b}$$\{e^{b}_{n}\}_{n=1}^{N}$ object query $m^{f}$$\{m^{f}_{n}\}_{n=1}^{N}$ binary mask
$z$ $z=\{(p_{n},e_{n},m_{n})\}_{n=1}^{N}$ prediction outputs
@@ 5.1 Datasets & Implementation
GraSP (GRASP). The GraSP dataset [4] is a new curated benchmark that models surgical scene understanding. This dataset contains 13 sequences, officially divided into 8 for training and 5 for testing. This dataset provides multi-granularity including short-term segmentation and long-term recognition annotations. Each frame has a high resolution of 800$\times$1280. It’s noteworthy that GraSP provides monocular consecutive frames but only pixel-wise segmentation annotation for 3449 sparse frames. We use pre-defined training and validation splits from [4].
@@ 5.3 Ablation Study
Inference Time. We analyze the inference time of each proposed component as Table 8, where “Num” means clip length. In comparison with baseline model, LOCOSTE requires additional computation for temporal and stereo contexts, but the performance gain over additional latency is substantial. Moreover, LOCOSTE can unplug the high-latency LACls module to make trade-off in terms of both latency and performance. Decreasing the duration of stereo clips and incorporating memory mechanisms can also enhance inference efficiency. We expect that inference optimization methods (e.g., TensorRT) can further reduce the latency of large model for long perception range settings.
@@
1 0.22 75.12 71.68 44.48
1 0.33 78.34 74.13 47.88
ST 4 0.51 81.54 78.16 57.86
ST 8 1.01 81.89 78.52 61.26
ST 4 0.98 81.91 78.11 57.95
ST 8 1.47 82.31 78.56 61.21
@@ 5.4 Qualitative Analysis
Fig. 10: Temporal consistency with LACOSTE and some video object segmentation(VOS).
@@
Table 10: VOS performance of LACOSTE and some VOS methods.
Method VOS Metrics
@@
XMEM 65.0 72.2 68.6
STM 60.0 65.7 62.9
STCN 63.2 68.2 65.7
LACOSTE(B) 88.1 87.8 88.0
Temporal Consistency of Segmentation Results. We further analysis temporal consistency of segmentation results and make comparison with some video object segmentation (VOS) methods including space-time memory network [35], STCN [10] and XMEM [9]. Given the absence of video annotations for most surgical datasets, we employed pretrained Video Object Segmentation (VOS) models for inference. Specifically, utilizing the segmentation annotation from the initial frame, VOS methods, leveraging the officially pretrained parameters, infer masks across a sequence of eight frames. For the LACOSTE model, we maintained the original inference process without utilizing the initial frame annotation. The visualization results are depicted in Figure 10. Additionally, we present several VOS evaluations, including Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J}\&\mathcal{F}$, to assess temporal consistency in Table 10. Both qualitative and quantitative analyses on examples indicate that LACOSTE achieves comparable temporal consistency. It is important to clarify that our intention is not to assert a superior temporal consistency compared to VOS methods. Instead, we aim to offer researchers potential insights and alternatives for establishing temporal correlations in scenarios where video-level annotations are unavailable.

### v3

@@ 1 Introduction
For SIS, the effectiveness of query-based segmentation (QBS) methods has been validated by previous works. The findings of [6] indicate that instrument misclassification primarily leads to low performance of current SIS methods. While most methods yield satisfactory results in terms of both the bounding box and segmentation mask, they frequently misclassify the output box or mask. Therefore, we follow the QBS paradigm and derive our method, grounded on three limitations of current methods, to mitigate instrument misclassification for surgical scenario. Firstly, surgical frames are recorded by stereo cameras and have inherent stereo attributes that are largely ignored by previous works. The complementary depth-aware information from stereo frame can enhance current view features, which help instrument localization and recognition for complex scenes. From this, we design a disparity-guided feature propagation (DFP) module, which integrates features from stereo views with an offline disparity estimation network explicitly. DFP can be inserted into QBS baseline without additional trainable parameters. To generalize well to a monocular setting, we also propose a pseudo stereo mechanism, which generates complementary right frames within multiple disparity ranges. Some examples are illstrated in Figure 3.
@@ 2 Related Work
$m$$\{m_{n}\}_{n=1}^{N}$ binary mask $p^{b}$$\{p^{b}_{n}\}_{n=1}^{N}$ class probability
$e$$\{e_{n}\}_{n=1}^{N}$ object query $m^{b}$$\{m^{b}_{n}\}_{n=1}^{N}$ binary mask Final Output
$\mathcal{Q}$ $\mathcal{Q}=\{q_{n}\}_{n=1}^{N}$ initial query $e^{b}$$\{e^{b}_{n}\}_{n=1}^{N}$ object query $p^{f}$$\{p^{f}_{n}\}_{n=1}^{N}$ class probability
$z$ $z=\{(p_{n},e_{n},m_{n})\}_{n=1}^{N}$ prediction outputs $m^{f}$$\{m^{f}_{n}\}_{n=1}^{N}$ binary mask
@@ 5.1 Datasets & Implementation
GraSP (GRASP). The GraSP dataset [5, 4] is a new curated benchmark that models surgical scene understanding. This dataset contains 13 sequences, officially divided into 8 for training and 5 for testing. This dataset provides multi-granularity including short-term segmentation and long-term recognition annotations. Each frame has a high resolution of 800$\times$1280. It’s noteworthy that GraSP provides monocular consecutive frames but only pixel-wise segmentation annotation for 3449 sparse frames. We use pre-defined training and validation splits from [4].
@@ 5.3 Ablation Study
0.219 75.12 71.68 44.48
0.323 78.34 74.13 47.88
ST 0.343 81.89 78.52 61.26
ST 0.371 82.31 78.56 61.21
Inference Time. We analyze the inference time of each proposed component as Table 8. For improving model efficiency and reducing inference time, we introduce a memory bank to record intermediate results including features, embeddings and predictions which can be reused across different clips. Furthermore, we also record corresponding valid object queries indices which can reduce the computations of LACls. In comparison with baseline model, LACOSTE requires additional computation for temporal and stereo contexts, but the performance gain over additional latency is substantial. We expect that inference optimization methods (e.g., TensorRT) can further reduce the latency of model.
@@ 5.4 Qualitative Analysis
Table 10: Evaluation on temporal consistency of segmentation results.
Method EV18
@@
XMEM 70.2 72.3 72.0 72.2
XMEM(F) 72.6 73.1 73.6 73.4
SAM2 82.6 81.4 81.5 81.4
SurgSAM-2 84.4 84.3 84.0 84.1
SurgSAM-2(P) 84.1 84.2 83.8 84.0
LACOSTE(S) 85.2 85.8 85.3 85.5
Temporal Consistency of Segmentation Results. We further analyze our temporal consistency of segmentation results together with some video object segmentation (VOS) methods including SAM2 [36, 38] and SurgSAM-2 [28] which is based on SAM2 and fine-tuned with EV18. We also supplement fine-tuned results of XMEM [10] with the same dataset, represented as XMEM(F). Actually, ours and VOS paradigms are inherently tailored for distinct tasks with different challenges. It is important to clarify that our intention is not to assert a superior temporal consistency compared to VOS methods. Instead, we aim to offer researchers potential insights and feasible alternatives for establishing temporal correlations in scenarios where video-level annotations are unavailable. The temporal consistency of segmentation results, which is not the primary focus or contribution of this paper, is an inherited outcome by mitigating misclassification and improving segmentation effectiveness. Particularly, VOS methods receive the annotated first frame for each object of the video sequence and propagate them through time dimension. For the LACOSTE model, we maintain the original inference process without utilizing any annotation. We present several VOS metrics, including Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J}\&\mathcal{F}$, together with semantic segmentation metric Ch_IoU in Table 10. We use category IDs as object IDs in these experiments. For all metrics, LACOSTE achieves comparable or favorable results with the other methods. In fact, our method does not require any annotations during inference process while the competing methods require so. Additionally, we substitute the annotations of reference frames with our predictions for SurgSAM-2 to validate the cooperative potential, denoted as SurgSAM-2(P). Our approach can deliver semantic prompts for category-agnostic SAM2-based VOS methods when reference annotations are unavailable.